THE AILES IR AND THZ HIGH RESOLUTION SPECTROSCOPY BEAMLINE AT SOLEIL : CHARAC-TERISTICS AND PERFORMANCES

PASCALE ROY, JEAN-BLAISE BRUBACH, MATHIEU ROUZIÈRES, MICHEL VERVLOET, OLIVIER PIRALI^{a}, DIDIER BALCON^{a}, FRIDOLIN KWABIA-TCHANA^{b}, \underline{LAURENT MANCERON}^{c}, Synchrotron SOLEIL, L’Orme des Merisiers, Saint-Aubin - BP 48, 91192 Gif-sur-Yvette CEDEX, France.

A new infrared beamline (Beamline AILES) is now operating at the third generation Synchrotron Radiation source SOLEIL near Saclay, France. This beamline makes use of infrared synchrotron radiation from both edge emission and constant field emission of a bending magnet, with optics optimised for extracting a large solid angle (20 × 78 mrad^{2}) and thus for operation in the far infrared region and use of high resolution (0.001 cm^{−1}) with high flux and brilliance. The expected and measured performances in term of flux, spatial distribution of the photons, spectral range and stability are discussed. Comparison with results obtained with standard laboratory sources show a 5 to 12-fold enhancement in signal-to-noise ratio over the 50 to 500 cm^{−1} range, for high resolution measurements. The optical system, spectroscopic stations and workspace are described, as well as the beamtime application procedures.

aLaboratoire de Photophysique Moléculaire, Université Paris-Sud, 91405 Orsay Cedex, France

bLab. Interuniversitaire des Systèmes Atmosphériques, CNRS-Paris 7 and 12, UMR 7583, 61 av. du Général de Gaulle, Créteil France

cLab. Dynamique, Interactions et Réactivité, CNRS-UPMC, UMR 7075, case 49, 4 place Jussieu, 75252 Paris Cedex, France